# Planar Collisionless Shock Simulations with Semi-Implicit Particle-in-Cell Model FLEKS

### v1

@@ 2.2 Local 2D Simulations
The simulated quasi-parallel shock is not steady nor in an equilibrium state, with the shock front moving gradually along the x-direction. We next check the particle behaviors at a fixed location to account for the spatiotemporal evolution. Figure 7 displays the time evolution of ion and electron distributions within a $5\,d_{i,\text{up}}\times 5\,d_{i,\text{up}}$ box centered at $x=-62.5\,d_{i.\text{up}}$. This location, initially downstream of the shock, transitions to the upstream region as the shock front propagates in the -x direction. Each column represents a snapshot in time, while each row shows the velocity space density projections onto the $V_{x}-V_{y}$ and $V_{x}-V_{z}$ planes for ions and electrons, respectively. At $t=0$, the plasma is initialized with single Maxwellian distributions based on MHD solutions. By $t=7.6\,\omega_{ci,\text{up}}^{-1}$, the transition to a kinetic state is evident, with the emergence of a non-Maxwellian but nearly isotropic ion distribution. At $t=13.3\,\omega_{ci,\text{up}}^{-1}$, the shock front passes by, shown by the central drift path in Figure7c-d. A distinct ion beam propagating upstream forms by $t=15.2\,\omega_{ci,\text{up}}^{-1}$, centered at $V_{ibx}\simeq 3\,V_{Ai,\text{up}}$. This field-aligned beam, with a relative velocity of $V_{0}=V_{0b}-V_{0m}\simeq 10\,V_{Ai,\text{up}}$ with respect to the upstream Maxwellian population, is a characteristic feature of the foreshock region. However, this beam structure is transient and also affected by the local magnetic field orientation (Figure 7g-h). By $t=19\,\omega_{ci,\text{up}}^{-1}$, it has dissipated, leaving a more diffuse and sparse upstream-moving population afterwards (Figure 7k-l). As the shock front continues to propagate in the -x direction, this fixed spatial location near $x=-62.5\,d_{i,\text{up}}$ samples the diffuse ion population that was previously further upstream (Figure 7g-h). The phase space densities of this diffuse population are typically below 1% of the peak density in the upstream core population. By the final snapshot at $t=20\,\omega_{ci,\text{up}}^{-1}$ (Figure 7i-j), the diffuse ion population becomes even more sparse. This time evolution complements the spatial comparisons of ion distribution in Figure 6k-t, providing a comprehensive picture of the spatiotemporal dynamics across the quasi-parallel shock.
@@
The lower half of Figure 7 shows the corresponding time evolution of the electron velocity distribution. As the shock front approaches (Figure 7m-p), the electron temperature increases while the distribution remains Maxwellian. At later times $t>15\,\omega_{ci,\text{up}}^{-1}$, when the observation location moves into the foreshock region, the electron distribution becomes non-Maxwellian with diffuse reflected species. This non-equilibrium distribution indicates a source of free energy that can potentially drive the small-scale waves observed in conjunction with the large-scale ULF magnetosonic waves. These waves, if exist, are likely triggered by sub-ion-scale instabilities, distinct from the ion-ion instability responsible for the magnetosonic waves [S. P. Gary, 1993]. However, in this planar shock setup with low mass ratio $m_{i}/m_{e}=25$, the sub-ion scale waves are not obvious.
@@
To track the time evolution of reflected ions, we again look at the $x-V_{ix}$ ion phase space distributions presented in Figure 8 from $t=0$ to $19\,d_{i,\text{up}}$. The whole simulation domain in the x-direction from $-150\,d_{i,\text{up}}$ to $250\,d_{i,\text{up}}$ is shown to facilitate a detailed analysis of the foreshock region. The initial MHD state (Figure 8a) contains no reflected ions before encountering the shock. As the kinetic system evolves, reflected ions appear early in the simulation (Figure 8b) and rapidly populate the upstream region (Figure 8c). However, the distinct beam-like structures extend only up to $\lesssim 10\,d_{i,\text{up}}$ ahead of the shock and are not present in all snapshots. Intriguingly, the growth of ULF magnetosonic waves is observed to commence after the formation of reflected ion species [L. Turc et al., 2018, 2025]. During the temporal delay between these snapshots, the initially coherent upstream traveling beam undergoes a transition to a diffuse population, a process that correlates with the saturation of wave growth. In parallel, the core species of the inflow are subject to subtle modulations resulting from the ULF wave perturbations.
Figure 9: Colored contours of ion density $n_{i}$ (a, d, g), velocity $U_{ix}$ (b, e, h), and ram pressure $P_{\text{ram}}$ (c, f, h) normalized to the upstream values close to the 2D $\theta_{Bn}=30^{\circ}$ quasi-parallel shock front at three consequent snapshots. The white lines in (a) are $B_{x}-B_{y}$ magnetic fields, and the gray quivers in (b) are $U_{ix}-U_{iy}$ ion velocities.
Finally we investigate the bulk variables to check the structures propagating through the quasi-parallel shock. Figure 9 presents a close-up view of the normalized ion density $n_{i}$, velocity $U_{ix}$, and ram pressure $P_{\text{ram}}$ near the shock front, spanning $x\in[-120,-40]\,d_{i,\text{up}}$ and $y\in[-70,70]\,d_{i,\text{up}}$, at three consequent snapshots $t=15.7,16.5,17.2\,\omega_{ci,\text{up}}^{-1}$. White lines with arrows and gray quivers represent in-plane magnetic field lines and ion velocities, respectively. Surface ripples are clearly visible near $x=-70\,d_{i,up}$ across this extended tangential domain. Several regions, both upstream and downstream, exhibit localized enhancements in ion density and velocity, leading to increased ram pressure $P_{\text{ram}}=m_{i}n_{i}U_{i}^{2}/2$. As shown by the time series, these localized structures are associated with upstream ULF wave perturbations and the propagation of nonlinear shocklets across the shock front into the downstream region.
@@
The localized enhancements in ram pressure are characteristic of jets [F. Plaschke et al., 2018], frequently observed in the terrestrial magnetosheath downstream of quasi-parallel shocks. In our simulation, across a tangential width of $140\,d_{i,\text{up}}\simeq 18,000\,\text{km}$, we find an occurrence rate of approximately one jet per $2\pi/\omega_{ci,\text{up}}$, or roughly 4 jets per minute, based on the inspecting time interval from $t=9$ to $20\,\omega_{ci}^{-1}$. Despite the lack of a clear definition for jets, our simulations reveal a jet occurrence rate higher than observational estimates of a few jets per hour. This discrepancy might be partly attributed to the generally smaller scale of our simulated jets ($\lesssim 10\,d_{i,\text{up}}$) compared to observed terrestrial magnetosheath jets, making them more challenging for satellites to detect. Furthermore, in our 2D planar shock simulations, the jets tend to dissipate rapidly in the downstream region, limiting their penetration depth to approximately $20\,d_{i,\text{up}}$. This limited penetration is likely influenced by the downstream oblique magnetic field orientation and the inherent constraint of the two-dimensional setup, which lacks the complexity of the third spatial dimension. Ultimately, these fundamental differences between our simplified 2D planar shock model and the actual global 3D magnetospheric configuration offer plausible explanations for the observed model-observation mismatch.
@@ 3 Discussion
Through decades of local planar kinetic simulations, we have already accumulated decent understandings of micro-scale dynamics in collisionless shocks. FLEKS was created with the goal of embedding local kinetic processes into a global MHD system, but it was not applied to study kinetic shock physics in the past due to the computational constraints and numerical difficulties. Here we demonstrate that FLEKS, with an improved field solver, can reveal the key kinetic processes and ULF waves across planar shocks at different shock-normal angles, using the average solar wind conditions at 1 au from WIND measurement. We notice that for Earth-like bow shocks with $\beta\simeq 1$ the ion phase space holes are less obvious than the low-$\beta$ cases as shown by many previous local full-PIC shock simulations, $\beta_{i}\simeq\beta_{e}\simeq 0.1$ (e.g. T. Umeda & R. Yamazaki [2006]; M. Nakanotani et al. [2022]). This can be easily understood as the larger thermal velocity extent smears out the holes in the velocity space, and different structures tend to overlap with each other. By experimenting with different initial shock states, we confirm that FLEKS also gives consistent results with previous local shock studies under different scenarios, for example the low Mach number subcritical quasi-perpendicular shock regime (MMS observation [D. B. Graham & Y. V. Khotyaintsev, 2025] and hybrid simulations [L. Ofman et al., 2009; L. Ofman & M. Gedalin, 2013] with periodic downstream structures under pressure equilibrium), and the low-$\beta$ supercritical shock regime [M. Scholer et al., 2003b] with the perpendicular shock reformation cycle. The comparisons clearly show that collisionless shock behaviors are very sensitive to the specific plasma regimes. These checks are useful for guiding the future global bow shock simulations under different upstream solar wind parameters, including but not limited to the average Mach number and plasma $\beta$ as we present above for validation, and also link back to the local studies of detailed shock physics.
@@
Early attempts to simulate mirror and EMIC waves often relied on simplified 1D hybrid simulations initialized with bi-Maxwellian ion distributions exhibiting high anisotropy ($P_{i\perp}/P_{i\parallel}=6$), as exemplified by C. P. Price et al. [1986]. These simulations, however, lacked a self-consistent treatment of the shock and its role in generating the anisotropy. Our experiments demonstrate that when the shock is included, the constraints imposed by reduced dimensionality become crucial. Specifically, 1D simulations are insufficient for capturing the downstream waves generated by perpendicular shocks. A 2D setup is the minimal requirement for accurately reproducing these anisotropy-driven waves.

### v2

@@ 3.2 Local 2D Simulations
The simulated quasi-parallel shock is not steady nor in an equilibrium state, with the shock front moving gradually along the x-direction. We next check the particle behaviors at a fixed location to account for the spatiotemporal evolution. Figure 7 displays the time evolution of ion and electron distributions within a $5\,d_{i,\text{up}}\times 5\,d_{i,\text{up}}$ box centered at $x=-62.5\,d_{i.\text{up}}$. This location, initially downstream of the shock, transitions to the upstream region as the shock front propagates in the -x direction. Each column represents a snapshot in time, while each row shows the velocity space density projections onto the $V_{x}-V_{y}$ and $V_{x}-V_{z}$ planes for ions and electrons, respectively. At $t=0\,\omega_{ci,\text{up}}^{-1}$, the plasma is initialized with single Maxwellian distributions based on MHD solutions. By $t=7.6\,\omega_{ci,\text{up}}^{-1}$, the transition to a kinetic state is evident, with the emergence of a non-Maxwellian but nearly isotropic ion distribution. At $t=13.3\,\omega_{ci,\text{up}}^{-1}$, the shock front passes by, shown by the central drift path in Figure7c-d. A distinct ion beam propagating upstream forms by $t=15.2\,\omega_{ci,\text{up}}^{-1}$, centered at $V_{ibx}\simeq 3\,V_{Ai,\text{up}}$. This field-aligned beam, with a relative velocity of $V_{0}=V_{0b}-V_{0m}\simeq 10\,V_{Ai,\text{up}}$ with respect to the upstream Maxwellian population, is a characteristic feature of the foreshock region. However, this beam structure is transient and also affected by the local magnetic field orientation (Figure 7g-h). By $t=19\,\omega_{ci,\text{up}}^{-1}$, it has dissipated, leaving a more diffuse and sparse upstream-moving population afterwards (Figure 7k-l). As the shock front continues to propagate in the -x direction, this fixed spatial location near $x=-62.5\,d_{i,\text{up}}$ samples the diffuse ion population that was previously further upstream (Figure 7g-h). The phase space densities of this diffuse population are typically below 1% of the peak density in the upstream core population. By the final snapshot at $t=20\,\omega_{ci,\text{up}}^{-1}$ (Figure 7i-j), the diffuse ion population becomes even more sparse. This time evolution complements the spatial comparisons of ion distribution in Figure 6k-t, providing a comprehensive picture of the spatiotemporal dynamics across the quasi-parallel shock.
@@
The lower half of Figure 7 shows the corresponding time evolution of the electron velocity distribution. As the shock front approaches (Figure 7m-p), the electron temperature increases while the distribution remains Maxwellian. At later times $t>15\,\omega_{ci,\text{up}}^{-1}$, when the observation location moves into the foreshock region, the electron distribution becomes non-Maxwellian with diffuse reflected species. This non-equilibrium distribution indicates a source of free energy that can potentially drive the small-scale waves observed in conjunction with the large-scale ULF magnetosonic waves. These waves, if exist, are likely triggered by sub-ion-scale instabilities, distinct from the ion-ion instability responsible for the magnetosonic waves [S. P. Gary, 1993]. However, in this planar shock setup with low mass ratio $m_{i}/m_{e}=25$ and $\Delta x=1\,d_{e}$, the sub-ion scale waves are not obvious.
@@
To track the time evolution of reflected ions, we again look at the $x-V_{ix}$ ion phase space distributions presented in Figure 8 from $t=0$ to $19\,d_{i,\text{up}}$. The whole simulation domain in the x-direction from $150\,d_{i,\text{up}}$ to $250\,d_{i,\text{up}}$ is shown to facilitate a detailed analysis of the foreshock region. The initial MHD state (Figure 8a) contains no reflected ions before encountering the shock. As the kinetic system evolves, reflected ions appear early in the simulation (Figure 8b) and rapidly populate the upstream region (Figure 8c). However, the distinct beam-like structures extend only up to $\lesssim 10\,d_{i,\text{up}}$ ahead of the shock and are not present in all snapshots. Intriguingly, the growth of ULF magnetosonic waves is observed to commence after the formation of reflected ion species [L. Turc et al., 2018, 2025]. During the temporal delay between these snapshots, the initially coherent upstream traveling beam undergoes a transition to a diffuse population, a process that correlates with the saturation of wave growth. In parallel, the core species of the inflow are subject to subtle modulations resulting from the ULF wave perturbations.
Figure 9: Colored contours of ion density $n_{i}$ (a, d, g), velocity $U_{ix}$ (b, e, h), and ram pressure $P_{\text{ram}}$ (c, f, h) normalized to the upstream values close to the 2D $\theta_{Bn}=30^{\circ},\mathrm{M}_{\mathrm{A}}=12$ quasi-parallel shock front at three consequent snapshots. The white lines in (a) are $B_{x}-B_{y}$ magnetic fields, and the gray quivers in (b) are $U_{ix}-U_{iy}$ ion velocities.
Finally we investigate the plasma moments to check the structures propagating through the quasi-parallel shock. Figure 9 presents a close-up view of the normalized ion density $n_{i}$, velocity $U_{ix}$, and ram pressure $P_{\text{ram}}$ near the shock front, spanning $x\in[-120,-40]\,d_{i,\text{up}}$ and $y\in[-70,70]\,d_{i,\text{up}}$, at three consequent snapshots $t=15.7,16.5,17.2\,\omega_{ci,\text{up}}^{-1}$. White lines with arrows and gray quivers represent in-plane magnetic field lines and ion velocities, respectively. Surface ripples are clearly visible near $x=-70\,d_{i,up}$ across this extended tangential domain. Several regions, both upstream and downstream, exhibit localized enhancements in ion density and velocity, leading to increased ram pressure $P_{\text{ram}}=m_{i}n_{i}U_{i}^{2}/2$. As shown by the time series, these localized structures are associated with upstream ULF wave perturbations and the propagation of nonlinear shocklets across the shock front into the downstream region.
@@
The localized enhancements in ram pressure are characteristic of jets. In run 14, across a tangential width of $140\,d_{i,\text{up}}\simeq 18,000\,\text{km}$, we find an occurrence rate of approximately one jet per $2\pi/\omega_{ci,\text{up}}$, or roughly 4 jets per minute, based on the inspecting time interval from $t=9$ to $20\,\omega_{ci,\text{up}}^{-1}$. The jets in the 2D planar shock simulations tend to dissipate rapidly in the downstream region, limiting their penetration depth to approximately $20\,d_{i,\text{up}}$.
3.2.3 Downstream Magnetic Reconnection and Jet Dynamics in Quasi-Parallel Shocks
Figure 10: Colored contours of out-of-plane current density $J_{z}$, electron velocity $U_{ey}$, and nongyrotropy measures $\sqrt{Q_{e}}$ and $D_{ng,e}$ in the 2D $\theta_{Bn}=30^{\circ}$ quasi-parallel shock run 14 ($\mathrm{M}_{\mathrm{A}}=7$, column a-c) and 15 ($\mathrm{M}_{\mathrm{A}}=12$, column d-f) at three consecutive snapshots, overplotted with in-plane magnetic field lines. Reconnection X-points and O-points are marked with cyan crosses and green circles, respectively.
The turbulent nature of the quasi-parallel shocks and the propagation of foreshock waves make the shock downstream region an ideal location for triggering magnetic reconnection [S. Wang et al., 2024] and forming jets [J. Ren et al., 2023]. To resolve key kinetic processes downstream of the shock, we further perform a comparative analysis of three 2D quasi-parallel shock simulations (runs 14, 15, and 16), focusing on magnetic reconnection and jet dynamics. We select $\sqrt{Q_{e}}$ and $D_{\text{ng,e}}$ as representatives from the list of electron non-gyrotropy measures that are enhanced at the reconnection sites [H. Zhou et al., 2020]. Figure 10 presents close-up views of the normalized out-of-plane current density $J_{z}$, electron velocity $U_{ey}$, and nongyrotropy measures $\sqrt{Q_{e}}$ and $D_{\text{ng,e}}$ downstream of the shock front, spanning $45\,d_{i,\text{up}}$ in x and $40\,d_{i,\text{up}}$ in y, at three consecutive snapshots from runs 14 and 15, which differ only in the upstream Mach number, overplotted with in-plane magnetic field lines. We identify the reconnection X and O points by finding the local saddle/extrema from the 2D magnetic flux function $\psi$, defined such that $\mathbf{B}=\nabla\psi\times\hat{z}$ (i.e. $B_{x}=\partial\psi/\partial y$ and $B_{y}=-\partial\psi/\partial x$).
In both cases, we observe magnetic reconnection occurring downstream of the shock front. This process is marked by the formation of current sheets (first row in Figure 10), which subsequently break up to form magnetic X-points (reconnection sites) and O-points (plasmoids). Current sheets form intermittently downstream of the shock front, which can be directly linked to the steepening and compression of foreshock waves as they are convected through the shock transition. During the shock reformation process at $t=7.2\,\omega_{ci,\text{up}}^{-1}$ in Run 15, we find a bunch of small reconnection sites and plasmoids between the previous shock front and the newly forming shock front (Figure 10f). Comparing Run 14 ($\text{M}_{\text{A}}=7$) and Run 15 ($\text{M}_{\text{A}}=12$) reveals a clear dependence on the Mach number. The higher Mach number in Run 15 leads to a larger out-of-plane current density $J_{z}$, more strongly curved magnetic field lines, and a significantly higher number of local reconnection X/O points. The plasmoids generated in the reconnection outflow have a length scale of up to $5\,d_{i,\text{up}}$.
An analysis of electron kinetic features, such as the nongyrotropy measure $\sqrt{Q_{e}}$ (M. Swisdak [2016], third row in Figure 10) and $D_{ng,e}$ (N. Aunai et al. [2013], forth row in Figure 10), shows that these quantities are enhanced not only near reconnection separatrices but also at the main shock front itself. This indicates that in the turbulent shock environment, other kinetic processes—such as wave-particle interactions at the shock front and instabilities within the compressed plasma—also contribute significantly to the generation of nongyrotropy.
Figure 11: Colored contours of out-of-plane current density $J_{z}$, ram pressure $P_{\text{ram}}$, ion density $n_{i}$, magnetic fields and electron and ion anisotropies in the 2D $\theta_{Bn}=0^{\circ},\mathrm{M}_{\mathrm{A}}=7$ quasi-parallel shock run 16 at four consecutive snapshots, overplotted with in-plane magnetic field lines. Reconnection X-points and O-points are marked with cyan crosses and green circles, respectively.
To further explore these dynamics in a different geometry, we present various quantities from the $\theta_{Bn}=0^{\circ}$ parallel shock (run 16) in Figure 11. We zoom-in to a local region that covers about $50\,d_{i,\text{up}}$ upstream/downstream from the shock front and $80\,d_{i,\text{up}}$ wide in the shock tangential direction. Magnetic field lines are shown in black and white on top of the colored variables. This run shares the same Alfvén Mach number as run 14. We find many more local reconnection X (cyan crosses) and O (green circles) points in the fully-parallel case, which are possibly linked to the more turbulent current sheet formation and reconnection processes in the shock transition layer for the small cone angle case. The out-of-plane current density $J_{z}$ also generates more curved magnetic fields.
The second row in Figure 11 shows the ram pressure $P_{\text{ram}}$, highlighting jet structures propagating downstream across the shock front, which dissipates over time. The jets in this full PIC simulation can penetrate up to $20\,d_{i,\text{up}}$ into the downstream region before dissipation. We also observe jets with bow-wave-like structures at the front (Figure 11 a2-c2 near $x=-40\,d_{i,\text{up}}$). In the upstream of the shock, we can see clear transverse magnetic oscillations from the magnetosonic-whistler modes (Figure 11 row 5-6). At a later time $t=15.2\,\omega_{ci,\text{up}}^{-1}$, filaments of ion density $n_{i}$ along the magnetic field direction appear with length scale $\sim 2\,d_{i,\text{up}}$ in the transverse direction. These filaments are quasi-equilibrium structures after the saturation of Weibel instability [R. L. Morse & C. W. Nielson, 1971] with an anti-correlation between the dynamic pressure and magnetic pressure.
@@ 4 Discussion
Through decades of local planar kinetic simulations, we have already accumulated decent understandings of micro-scale dynamics in collisionless shocks. FLEKS was created with the goal of embedding local kinetic processes into a global MHD system, but it was not applied to study kinetic shock physics in the past due to the computational constraints and numerical difficulties. Here we demonstrate that FLEKS, with an improved field solver, can reveal the key kinetic processes and ULF waves across planar shocks at different shock-normal angles, using the average solar wind conditions at 1 AU from WIND measurement. We notice that for Earth-like bow shocks with $\beta\simeq 1$ the ion phase space holes are less obvious than the low-$\beta$ cases as shown by many previous local full-PIC shock simulations, $\beta_{i}\simeq\beta_{e}\simeq 0.1$ (e.g. T. Umeda & R. Yamazaki [2006]; M. Nakanotani et al. [2022]). This can be easily understood as the larger thermal velocity extent smears out the holes in the velocity space, and different structures tend to overlap with each other. By experimenting with different initial shock states, we confirm that FLEKS also gives consistent results with previous local shock studies under different scenarios, for example the low Mach number ($\text{M}_{\text{A}}\sim 3$) subcritical quasi-perpendicular shock regime (MMS observation [D. B. Graham & Y. V. Khotyaintsev, 2025] and hybrid simulations [L. Ofman et al., 2009; L. Ofman & M. Gedalin, 2013] with periodic downstream structures under pressure equilibrium), the low-$\beta$ supercritical shock regime [M. Scholer et al., 2003b] with the perpendicular shock reformation cycle, moderate Mach number ($\text{M}_{\text{A}}\sim 10$) supercritical parallel shock regime with downstream reconnection and jets [S. Wang et al., 2024; J. Ren et al., 2024], and high Mach number ($\text{M}_{\text{A}}\sim 60$) supercritical shocks with filamentary Weibel instability [R. L. Morse & C. W. Nielson, 1971] and non-resonant streaming instability (i.e. Bell instability [A. Bell, 2004]). The comparisons clearly show that collisionless shock behaviors are very sensitive to the specific plasma regimes. These checks are useful for guiding the future global bow shock simulations under different upstream solar wind parameters, including but not limited to the average Mach number and plasma $\beta$ as we present above for validation, and also link back to the local studies of detailed shock physics.
4.1 Dimensionality: 1D Artifacts and the Need for Higher Dimensions
@@
Early attempts to simulate mirror and EMIC waves often relied on simplified 1D hybrid simulations initialized with bi-Maxwellian ion distributions exhibiting high anisotropy ($P_{i\perp}/P_{i\parallel}=6$), as exemplified by C. P. Price et al. [1986]. These simulations, however, lacked a self-consistent treatment of the shock and its role in generating the anisotropy. Our experiments demonstrate that when the shock is included, the constraints imposed by reduced dimensionality become crucial. It is important to recognize that our 1D simulation (Figure 1) neglects parallel heating along the z-direction. Consequently, no proton thermalization occurs along $V_{iz}$ in the downstream and ramp regions (Figure 1j and 1n). This leads to highly anisotropic downstream ion distributions ($P_{i\perp}/P_{i\parallel}>8$), a clear artifact of the reduced dimensionality. As demonstrated in the 2D quasi-perpendicular simulation (Figure 5), downstream parallel and oblique propagating waves, absent in the 1D case, play a crucial role in redistributing free energy associated with temperature anisotropy. This finding demonstrates that 1D simulations are insufficient for capturing the downstream waves generated by perpendicular shocks. The 2D simulation, incorporating realistic solar wind conditions upstream of Earth [C. S. Salem et al., 2023], provides a more accurate representation of the terrestrial bow shock compared to the 1D case, underscoring the importance of at least 2D modeling for capturing the complex dynamics of quasi-perpendicular shocks, including the interplay of anisotropy-driven instabilities and wave generation.
It is also crucial to acknowledge certain limitations inherent in the 1D planar shock setup for quasi-parallel shocks. Firstly, with only one spatial degree of freedom, wave propagation is constrained to the x-direction. We established our background upstream magnetic field at a shock-normal angle $\theta_{Bn}=30^{\circ}$ relative to the -x direction. Consequently, this oblique wave angle remains fixed and may not align with the maximum wave growth direction predicted by linear theory. However, E. A. Lucek et al. [2002] reported from in-situ satellite measurements that the average $\theta_{Bn}$ is approximately $20^{\circ}$, a value inconsistent with the prediction for maximum parallel propagating linear wave growth. This discrepancy illustrates the complexity of wave behavior in real-world environments compared to idealized simulations. Secondly, the periodic condition applied in the y and z direction limits the growth of structures that are intrinsically high-dimensional, such as downstream magnetic reconnection, plasma jets and return flows, which are observed in our 2D simulations (Figure 9, 10 and 11).
4.2 Interpreting Downstream Structures in 1D
In our 1D quasi-perpendicular simulation, correlated periodic downstream fluctuations are observed (Figure 1a, c, d). These postshock oscillations are not indicative of transmitted or generated waves. Instead, they are periodic structures that arise from ion ring distributions. This is true despite their exhibiting density and magnetic field correlations similar to compressional fast magnetosonic waves [C. T. Russell et al., 2009]. Interpretations of these structures as damping solitons [R. Sagdeev, 1966; C. Kennel, 1967] are inconsistent with the shock’s $\beta$ and Mach number. These structures also differ from the stationary downstream oscillations analyzed in L. Ofman et al. [2009]; L. Ofman & M. Gedalin [2013], where subcritical shocks maintain total pressure balance with anti-correlated magnetic field and ion pressure variations. While their descriptions involve non-gyrotropic downstream ions, our simulations clearly show a symmetric velocity distribution in the perpendicular $V_{ix}-V_{iy}$ plane. The ramp width in our simulation aligns with observations from early AMPTE satellite measurements [S. N. Walker et al., 1999a, b] and subsequent Cluster and THEMIS data [Y. OHobara et al., 2010]. Furthermore, the revealed ion reflection is consistent with the model proposed in the Appendix of Y. V. Khotyaintsev et al. [2024], which falls into the category where ions, upon reflection, are decelerated by $\Delta\phi$ and then undergo cyclotron gyration due to the increased magnetic field in the ramp, leading to a broad distribution of the reflected population within the ramp. The inclusion of electron kinetics in FLEKS introduces electrostatic electron pressure modulation by ion and magnetic pressure variations (Figure 1c, h). This effect is absent in hybrid models that treat electrons as a massless fluid in an ideal thermodynamic process (e.g. adiabatic if $\gamma=5/3$; isothermal if $\gamma=1$).
4.3 Foreshock Wave Dynamics: Shocklets, SLAMS, and Whistlers
Our simulations of quasi-parallel shocks capture the nonlinear evolution of upstream ULF waves. As the fast magnetosonic waves are convected back towards the shock by the super-Alfvénic upstream flow, they evolve into shocklets [M. Hoppe et al., 1981; L. B. Wilson III, 2016], i.e. steepened flanks characterized by small-scale fluctuations ($\delta B/B_{0}<2$) that function as miniature shocks. Short, Large Amplitude Magnetic Structures (SLAMS, or magnetic pulsations, S. J. Schwartz et al. [1992]) share similarities with shocklets but are defined by larger magnetic field fluctuations ($\delta B/B_{0}>2$) and potential soliton-like behavior. Consequently, our nonlinearly steepened ULF wave signatures (Figure 2) align more closely with observed shocklets, where nonlinear evolution is driven by interactions with energetic particle pressure gradients, and the ensemble of these structures forms the shock transition (e.g., [N. Dubouloz & M. Scholer, 1995]). The downstream magnetic fluctuations also exhibit higher amplitudes and broader spatial extents, consistent with observational data from Earth’s quasi-parallel magnetosheath (e.g., E. A. Lucek et al. [2002]).
The presence of whistler precursors is sensitive to the mass ratio (Figure 3). We observe clear whistler waves upstream of the shock front with wavelength $\lambda_{w}\sim 2d_{i,\text{up}}$ in larger mass ratio cases ($m_{i}/m_{e}=100,400$), where the whistler critical Mach number $\text{M}_{\text{w}}\gtrsim\text{M}_{\text{A}}$, consistent with the formation criterion (Equation 1). These whistler waves consistently appear on the leading edges in the first few ULF wavelengths ahead of the shock, which may be linked directly to the main shock front or the upstream steepened shocklets [L. B. Wilson III, 2016]. Similar whistler precursor occurrences are reported in full PIC simulations, e.g. K. Tsubouchi & B. Lembège [2004]; M. Nakanotani et al. [2022].
4.4 Downstream Jets and Reconnection
Our 2D quasi-parallel simulations show localized enhancements in ram pressure characteristic of jets [F. Plaschke et al., 2018], which are frequently observed in the terrestrial magnetosheath. In run 14 (Figure 9), we find an occurrence rate of approximately one jet per $2\pi/\omega_{ci,\text{up}}$, or roughly 4 jets per minute. Despite the lack of a clear definition for jets, our simulations reveal a jet occurrence rate higher than observational estimates of a few jets per hour. This discrepancy might be partly attributed to the generally smaller scale of our simulated jets ($\lesssim 10\,d_{i,\text{up}}$) compared to observed terrestrial magnetosheath jets, making them more challenging for sparse satellites to detect. The observed return flow (Figure 9b) is similar to satellite observation of sunward flows in the downstream magnetosheath (e.g., M. Archer et al. [2014]).
Furthermore, our simulations show that these jets dissipate rapidly. In the parallel shock case (Figure 11), the jets penetrate up to $20\,d_{i,\text{up}}$ before dissipation. We also observe bow-wave-like structures at the jet front (Figure 11 a2-c2), consistent with the hybrid simulations of J. Ren et al. [2024]. However, the jet scale sizes are shorter than the $\sim 100\,d_{i,\text{up}}$ penetration depth reported in J. Ren et al. [2024]. Part of this difference arises from the inclusion of kinetic physics across the electron and ion scales in the FLEKS model. Subion-scale micro-instabilities, which are absent in hybrid simulations where electrons are treated as a fluid, may be excited by the strong pressure anisotropies (Figure 11 row 7-8) and inhomogeneous flows near the shock front. These instabilities could provide an additional dissipation channel that hinders the deep penetration of jets, indicating a key physical process that requires a full kinetic treatment to be captured. This limited penetration in our 2D planar shock simulations, along with the inherent constraint of the two-dimensional setup, offers plausible explanations for the observed model-observation mismatch in jet occurrence and scale. Finally, our analysis of X/O points and electron nongyrotropy measures (Figure 10 row 3-4) confirms the presence of shocked reconnection [S. Wang et al., 2024] but also shows that in this turbulent environment, other kinetic processes also contribute significantly to nongyrotropy, making it a challenging metric for uniquely identifying reconnection sites.
4.5 Electron-to-Ion Temperature Ratio